OpenStudy (anonymous):

Express as a single fraction cos 105

OpenStudy (nurali):

Use the unit circle when you need exact values. 105° is not one of the values given on the unit circle, but 45° and 60° are. Since 105° = 45° + 60°, you can use the "angle addition formula" cos(α+β) = cos(α)×cos(β) - sin(α)×sin(β) cos(45°) = sin(45°) = √2/2 cos(60°) = 1/2 sin(60°) = √3/2 cos(105°) = cos(45°)×cos(60°) - sin(45°)×sin(60°) = (√2/2)(1/2) - (√2/2)(√3/2) = √2/4 - √2√3/4 = (√2 - √6)/4

OpenStudy (anonymous):

cos105=cos(90+15) => cos105=cos90cos15-sin90sin15 =>cos105=-sin15